ADDITIONS

Tuesday, 3.00 pm in 131 Mathias Jais: ” Inverse problems for parabolic partial differential equations with hysteresis” Wednesday 5.30 pm in 145: A. Boumenir: ”The determinant method for Sturm-Liouville Problems”. Friday 2.00 pm in 133 R. Sims ”Stability of the absolutely continuous spectrum of random Schrödingers operator on tree graphs.” Friday 2.30 pm in 133 S. Warzel ”Absolutely continuous spectrum of random quantum trees.” Friday 4.30 pm in 131: O.Ignatiev: ” On Krasovskii Criterion of Optimal Stabilization”

CANCELLATIONS

B.Belinskii: Tuesday, 3.00 pm in 131. M.Perel: Wednesday, 1.30 pm in 131. S.Warzel: Friday 2.00 pm in 133 (rescheduled, see Additions). R.Sims Friday 2.30 pm in 133 (rescheduled, see Additions). A.Boumenir: Friday 4.30 pm in 131 (rescheduled, see Additions). O.Ignatiev: Saturday 2.00 pm in 145 (rescheduled, see Additions).